Grade 12Mechanics

A gun moving at a speed of 30m/sec fires at an angle 30o with a velocity 150m/s relative to the gun. Find the distance between the gun and the projectile when projectile hits the ground. (g = 10 m/sec). how do we solve this question?

Kushi
11 Years agoGrade 12

3 Answers

akash
10 Years ago
velocity of projectile relative to ground =180 m/s now find the timde of flight. now relative displacement=urelative*time of flight 
Tushar Pagar
8 Years ago
 

Vertical component of velocity = 150 sin 30o  = 75 m/sec

Horizontal component of velocity relative to gun = 150 cos 30o = 75v3 m/sec

Horizontal component of velocity relative to ground = 75v3 + 30 ˜ 160m/sec

Time of flight   = (2\times75)/g = 15 s

Range of projectile = 160 × 15 = 2400 m

Distance moved by the gun and projectile = 2400 - 450 = 1950 m.

Abhijeet
5 Years ago
velocity of bullet wrt to gun comes out to be 180m/s.so total time of flight of bullet from projectile equation comes out to be 18 sec.range